OpenStudy (anonymous):

w=2wxyz solve for y pls explain steps to solve. Thx

OpenStudy (anonymous):

anyone know the answer??

OpenStudy (anonymous):

w = 2wxyz w/(2wxz) = 2wxyz/(2wxz) Divide both sides by 2wxz to isolate the y w/(2wxz) = y 1/2xz = y Cancel out the w since it appears in the numerator and denominator